OpenStudy (anonymous):

Find the lateral area, total area, and volume for the cylinder with the given measurement. r = 4, h = 5

OpenStudy (anonymous):

lateral area=2pirh=2*3.14*4*5=125.6 sq units total area=2pir^2+2pirh=2pir(r+h)=2*3.14*4(4+5)=226.08 sq units volume=pir^2h=3.14*16*5=251.2 cubic units